OpenStudy (shelby1290):

For the function g(x)=x^2-4x; simply: g(x+1)

OpenStudy (anonymous):

What type of math is this, shelby?

OpenStudy (shelby1290):

@rheaanderson341 it's functions

OpenStudy (dan815):

just stick x+1 where u see x.. and simplify

OpenStudy (dan815):

g(x+1)=(x+1)^2-4*(x+1) simplify (x+1)^2-4*(x+1)

OpenStudy (dan815):

(x+1)^2-4*(x+1) (x+1)^2= x^2+2x+1 4*(x+1) =4x+4 (x+1)^2-4*(x+1) =(x^2+2x+1) - (4x+4) =x^2+2x-4x+1-4 =x^2-2x-3

OpenStudy (shelby1290):

ahhhh okay i get it now. Thank you